Your credit card has a current balance of $4,965.20. This balance is accruing interest at a nominal rate of 24.0% compounded monthly. A) Assuming you didn’t spend any more on this card, what uniform end-of-month payments over the next 36 months would be required to reduce the loan balance to zero at the end of three years from today? B) Suppose you plan on spending an additional $1,000 at the end of next year (end of month 12) and $1,000 at the end of the following year (end of month 24) on this same card. What uniform end-of- month payments over the next 36 months would be required to pay off the existing debt along with the additional new purchases?
Answer #1

A) monthly payments over next 36 months will be used to pay off the loan. Let the payments be x.

$4,965.20 = x*PVAF(2%, 36 Periods)

4965.20 = x*25.489

$194.80

B) present value of debt to be repaid = 4,965.2+ 1,000*PVF(2%, 12 Periods) + 1,000*PVF(2%, 24 Periods)

= 4965.2 + 788.5+ 621

=$6,374.7

Let x be the annual payments

6,374.7 = x*PVAF(2%, 36 Periods)

X = $250.10